Be sure to answer all parts. Calculate the molar mass of the following substances: (a) Li2CO3 g/mol (b) KNO3 g/mol (c) Mg3N2 g/mol
Answer #1

Answer a)

Molar mass of Li = 7.0 g/mol

Molar mass of C = 12.0 g/mol

Molar mass of O = 16.0 g/mol

So molar mass of Li2CO3= (7.0 g/mol x 2) + (12.0 g/mol x 1) + (16.0 g/mol x 3) = 74.0 g/mol

Answer b)

Molar mass of K = 39.1 g/mol

Molar mass of N = 14.0 g/mol

Molar mass of O = 16.0 g/mol

So molar mass of KNO3 = (39.1 g/mol x 1) + (14.0 g/mol x 1) + (16.0 g/mol x 3) = 101.1 g/mol

Answer C)

Molar mass of Mg = 24.3 g/mol

Molar mass of N = 14.0 g/mol

So molar mass of Mg3N2= (24.3 g/mol x 3) + (14.0 g/mol x 2) = 100.9 g/mol